What is the number of yes/no vote combinations with 9 voters?
Answer by jim_thompson5910(35256)

There are 2 possible choices and 9 slots (to represent each person), so there are 2*2*2*2*2*2*2*2*2 = 2^9 = 512 different ways for the voters to vote yes/no.